OpenStudy (anonymous):

Factor the expression x^2 − 5x − 24.

OpenStudy (lieutenantgeneral):

These factors pair up and multiply to -24. 1*(-24) = -24 2*(-12) = -24 3*(-8) = -24 4*(-6) = -24 (-1)*(24) = -24 (-2)*(12) = -24 (-3)*(8) = -24 (-4)*(6) = -24

OpenStudy (anonymous):

(x+3)(x-8) thank you!